Fashion Outfitters Income Statement In 000s 2015 350,000 370,000 411,000 235,000 246,000 270,000 115,000 124,000 141,000 90,000 92,000 90,000 25,000 32,000 51,000 16,000 18,000 20,000 41,000 50,000 71,000 Results 2017 2016 Revenue Less: Cost of Goods Sold (COGS) Gross Profit Less: Selling, General & Administrative Operating Income Plus: Depreciation & Amortisation EBITDA EBIT Plus: Other Income (expense) Profit Before Tax Less: Taxes Net Profit for the Yea 25,000 32,000 51,000 9,000 19,000 16,000 44,000 51,000 67,000 5,000 26,000 29,000 19,000 25,000 38,000 Dividends Paid Ordinary Shareholders 9,500 12,500 19,000 Additional Information Share Price at Year End Shares Outstanding- Ordinary Net Cash Flow from Operations (000s) Effective Tax Rate 5.65 6.10 6.45 18,000 78,000 78,000 8,600 5,300 7,600 30.4% 31.5% 32.1%

in 000s 2017 95,000 129,100 134 183,700 179,600 156,1 46.800 49,700 56,1 54.400 57,200 63,1 2016 2015 h & Cash Equivalents Trade Receivabies Current Assets 3,000 4,500 482,900 420,100 415 Total Current Assets Plant & Equipment 118.300 146,300 154,1 13.500 3,500 1 25.300 29,300 46 Total Assets E640,000 609,200 628 41,500 44,400 37,500 40,200 43, 16.400 19,800 19, Term Loan Payable Current Liabilities 4.900 4,700 9, 100,300 109,100 Total Current Liabilities Term Loan Payable Liabilties 77 200 97,500 107 38.900 34,000 39 Taxes 17,800 20,600 25, 234,200 261,200 Total Liabilities Retained Earnings Shareholders Equity Total Shareholders Equity 147,000 137,500 125 405,800 348,000 335 Total Liabilities & Shareholders Equity 640.000 609,200 6628 258.900 210500 Prepare and critically analyse the key financial ratios for the the company over the three-year period, including at least one competitor, which must also operate internationally. Be sure to include relevant ratios from the four categories of financial analysis: liquidity, efficiency, profitability, and solvency (Marks: 30) Senior management of Fashion Outfitters have been discussing making an investment in a dedicated credit collections centre, which is expected to improve the companys collection period by at least 7 days Critically evaluate the impact this will have on the Cash Conversion Cycle and, subsequently, on the companys Working Capital requirements. (Do not include the valuation of the investment in your answer.) Marks: 20
Answer #1
Fashion Outfitters The Gap Inc.
2017 2016 2015 2017 2016 2015
Liqudity ratios
Quick ratio =(195000+183700+46800)/100300 =(129100+179600+49700)/109100 =(134000+156100+56100)/120900
=(cash+ST investment+ currenct receivable)/ current liabilities 4.24 3.29 2.86 1.01 0.83 1.09
Improving over 3 years and better than competition
Current ratio =482900/100300 =420100/109100 =415200/120900
=current assets/ current liabilities 4.81 3.85 3.43 1.76 1.57 1.93
Improving over 3 years and better than competition
Working capital =482900-100300 =420100-109100 =415200-120900
=current assets- current liabilities                                                 382,600                                                 311,000                                                 294,300 1,862,000 1,450,000 2,162,000
Improving over 3 years. Overall, Fashion is managing in capital efficiently
Solvency ratios
Debt to Equity ratio =234200/405800 =261200/348000 =293300/335500
=Total Liabilities/ Total equity 0.58 0.75 0.87 0.39 0.39 0.41
Improving over 3 years but worse than competition
Equity ratio =405800/640000 =348000/609200 =335500/628800
=Total Equity/ Total assets 0.63 0.57 0.53 0.84 0.81 0.83
Improving over 3 years but worse than competition
Debt ratio =234200/640000 =261200/609200 =293300/628800
=Total Liabilities/ Total assets 0.37 0.43 0.47 0.16 0.19 0.17
Improving over 3 years but worse than competition. Overall, Fashion has been improving its capital structure but it could do better
Efficiency ratios
Working capital =482900-100300 =420100-109100 =415200-120900
=current assets- current liabilities                                                 382,600                                                 311,000                                                 294,300 1,862,000 1,450,000 2,162,000
'Improving over 3 years.
Asset turnover =350000/((640000+60900)/2) =370000/((609200+628800)/2)
=net sales/ average total assets 1.00 0.60 2.03 1.98
'Improving over 2 years.
Inventory turnover =235000/((54400+57200)/2) =246000/((57200+63100)/2)
=COGS/ Average inventory 4.21 4.09 5.33 5.36
'Improving over 2 years.
Days sales in inventory =(54400/235000)*365 =(57200/246000)*365 =(63100/270000)*365
=ending inventory/COGS*365 84.49 84.87 85.30 67.63 67.84 67.96
Almost stable over the period but could be improved as compered to competition
Profitability ratios
Gross profit margin =115000/350000 =124000/370000 =141000/411000
=gross profit/ net sales 32.86% 33.51% 34.31% 38.46% 39.87% 36.44%
Decreasing over 3 years. Worse than competition
Net profit margin =19000/350000 =25000/370000 =38000/411000
=net profit/ net sales 5.43% 6.76% 9.25% 5.57% 6.37% 6.79%
Decreasing over 3 years. Worse than competition
Return on capital employed =25000/405800 =32000/348000 =51000/335500
=net operating profit/ employed capital 6.16% 9.20% 15.20% 22.76% 23.88% 32.80%
Decreasing over 3 years. Worse than competition
Return on equity =19000/405800 =25000/348000 =38000/335500
=net income/ total equity 4.68% 7.18% 11.33% 28.84% 22.76% 23.86%
Decreasing over 3 years. Overall, the business is not profitable for the shareholders. It needs to improve the business model to be more profitable
Cash conversion cycle
CCC = DIO + DSO - DPO =86.7+50.3-66.7 70.3
DIO = Average inventory/COGS per day =((54400+57200)/2)/235000*365 86.7
DSO = Average AR / Revenue per day =((46800+49700)/2)/350000*365 50.3
DPO = Average AP/COGS per day =((41500+44400)/2)/235000*365 66.7

If the collection period is improved by 7 days, the impact would be on DSO. The DSO would decrease to 43.3. This would lower the CCC to 63.3
The revenue per day is ~960,000. By improving AR by 7 days, the interes worth 960,000*7*Interedst rate/365 could be saved. This will improve the efficiency ratios as well